# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

### v1

@@ 9. Conclusion
In this paper, we identified five key findings in designers’ normal workflow with AI design tools. To address these existing gaps, we proposed PrototypeAgent, a novel multi-agent prototype generation system, delicately balancing automation with customization. Given human-provided descriptions and wireframe layouts, our system iteratively refines these inputs into engaging, high-fidelity design prototypes, maintaining aesthetic harmony and aligning with the design intent. Beyond generation, PrototypeAgent not only automates Implicit Intent Clarification for designers but also provides explainable intermediate results to enhance Human-Centered Intent-Design Alignment. Our quantitative and qualitative evaluations further corroborate the potential of our approach to significantly improve the UI / UX design process. Going forward, we will continue improving our work, like enabling automated wireframe generation and supporting dynamic component integration for a more universal, user-friendly, efficient and creative design process.

### v3

@@ 9. Conclusion
In this paper, we identified five key gaps in designers’ workflows with current AI-assisted design tools. To address these challenges, we introduced PrototypeFlow, a novel multi-module system that balances automation with customization. Given human-provided descriptions and wireframe layouts, our system iteratively refines these inputs into engaging, high-fidelity design prototypes, maintaining aesthetic harmony and aligning with the design intent. Beyond generation, PrototypeFlow not only automates design intent enhancement for designers but also provides editable intermediate results to enhance rapid regeneration. Our quantitative and qualitative evaluations further corroborate the potential of our approach to significantly improve the UI / UX design process. Going forward, we will continue improving our work, like enabling automated wireframe generation and supporting dynamic component integration for a more universal, user-friendly, efficient and creative design process.